OpenStudy (anonymous):

really confused! Morgan has nickels, dimes, and quarters in her piggy bank. In total, she has 45 coins, with a value of $5.25. if she has three more dimes than all the nickels and quarters combined, how many of each type of coin does she have?

OpenStudy (anonymous):

ok, let's call the nuckels n, dimes d, and quarters q. can you set up any systems of equations?

OpenStudy (anonymous):

n+d+q=45 5.25= 0.05n+0.10d+0.25q

OpenStudy (anonymous):

ok, there's one more you can make with the last sentence

OpenStudy (anonymous):

not quite sure what to do after that

OpenStudy (anonymous):

ok, wellshe has three more dimes that the total of nickels and quarters so d = 3 + n + q does that make sense?

OpenStudy (anonymous):

yes actually i got that far in class today, then i couldnt get past that

OpenStudy (anonymous):

ok, well substitute the third equation in for d into the first two equations

OpenStudy (anonymous):

5.25=0.05n+0.1(3+n+q)+0.25q but then you're still left with two variables

OpenStudy (anonymous):

do you isolate one of them and substitute again?

OpenStudy (anonymous):

exactly. solve those two equations as you would a noral system of equations

OpenStudy (anonymous):

oh so 45=n+(3+n+q)+q and 5.25=0.05n+0.1(3+n+q)+0.25q

OpenStudy (anonymous):

exactly! can you do it from here?

OpenStudy (anonymous):

yes thank you so much :)

OpenStudy (anonymous):

anytime :)

OpenStudy (anonymous):

Do you think I could get rid of the decimals and then put them in at the end? Do you think I would lose marks doing that?

OpenStudy (anonymous):

no, you wouldn't, but you have to be really careful. what you want to do is multiply the whole thing by 100, so that it's like you're doing it in terms of cents and not dollars, ok?

OpenStudy (anonymous):

Alright, I think I will still to decimals :P

OpenStudy (anonymous):

wait one last question

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Actually wrong one :P